OpenStudy (anonymous):

Verify the identity. 1-cos^2x=TanxCosxSinx

OpenStudy (anonymous):

got it?

mathslover (mathslover):

@sauravshakya direct answer dude.. please let the user do it by him/her selft

mathslover (mathslover):

*self

OpenStudy (anonymous):

SORRY.........

OpenStudy (anonymous):

MY mistake

mathslover (mathslover):

No problem but remember this from now on wards

OpenStudy (anonymous):

Don't forget he is an Ambassador..

OpenStudy (anonymous):

....yeahh

OpenStudy (anonymous):

1-cos^2 x =sin^2 x

mathslover (mathslover):

No no water.. He is my friend that is why I hoped that I had been polite rather been a rude person. ?

mathslover (mathslover):

*hope

OpenStudy (anonymous):

You took it seriously I am just kidding...

OpenStudy (anonymous):

okay yes I got that part. 1-cos^2x = sin^2x

OpenStudy (anonymous):

ya... we r friends..

OpenStudy (anonymous):

now, sin^2 = sinx * sinx

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

missing x on left..

OpenStudy (anonymous):

yeess

OpenStudy (anonymous):

oh ya.... thanx @waterineyes

OpenStudy (anonymous):

Now, sinx * sinx * cosx/cosx

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

He just multiplied and divide by cos(x)..

OpenStudy (anonymous):

where did you get the cosx/cosx from ?

OpenStudy (anonymous):

\[\sin(x) \cdot \sin(x) \times \frac{\cos(x)}{\cos(x)}\]

OpenStudy (callisto):

tanxsinxcosx = (sinx/cosx) sinxcosx = sinx * sinx * cosx / cosx

OpenStudy (callisto):

For the right side.

OpenStudy (anonymous):

@Callisto I am also thinking the same that now we should simplify right hand side to prove..

OpenStudy (cwrw238):

he used the identity tanx = sinx/cosx

OpenStudy (anonymous):

So, then I am leaving....

OpenStudy (anonymous):

Actually if we want to just prove by using left hand side then @sauravshakya is going right...

OpenStudy (callisto):

Uh-oh. I'm sorry....

OpenStudy (anonymous):

Carry on @sauravshakya

OpenStudy (anonymous):

okay so far I have sinx.sinx= sinx/cosx cosx.sinx now what do I do ?

OpenStudy (anonymous):

Use this : \[\frac{\sin(x)}{\cos(x)} = \tan(x)\]

OpenStudy (anonymous):

alright. so tanx=tanx right ?

OpenStudy (anonymous):

I mean to say that use this identity there...

OpenStudy (anonymous):

Check for colored part: \[\color{blue}{\frac{\sin(x)}{\cos(x)}} \cdot \sin(x) \cdot \cos(x)\]

OpenStudy (anonymous):

Can you replace that colored part ?

OpenStudy (anonymous):

u 1/cscx / 1/secx ?

OpenStudy (anonymous):

see you have: \[\color{blue}{\frac{\sin(x)}{\cos(x)}} \cdot \sin(x) \cdot \cos(x)\] And there is one formula which says that: \[\frac{\sin(x)}{\cos(x)} = \tan(x)\] Can you use this formula above ??

OpenStudy (anonymous):

umm no im kind of confused

OpenStudy (anonymous):

What is blue colored part can you write here ??

OpenStudy (anonymous):

Only write the part that I have shown you in blue color..

OpenStudy (anonymous):

sinx/cosx would be tanx

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

so just repalce it here: \[\color{blue}{\frac{\sin(x)}{\cos(x)}} \cdot \sin(x) \cdot \cos(x)\]

OpenStudy (anonymous):

only change the blue colored part with what you just said above..

OpenStudy (anonymous):

tanx.sinx.cosx. okay than what ?

OpenStudy (anonymous):

You want to do more in this ??

OpenStudy (anonymous):

do i have to or is that it?

OpenStudy (anonymous):

With your big eyeballs, see the right hand side of your question..

OpenStudy (anonymous):

LOL okay nvm thank you ! :)

OpenStudy (anonymous):

Ha ha ha.. Just kidding.. Welcome dear...

OpenStudy (anonymous):

Thanks to @sauravshakya and @Callisto too...